# Prompt Compression with Context-Aware Sentence Encoding for Fast and Improved LLM Inference

### v2

@@ 1 Introduction
The advent of large language models (LLMs) has triggered a surge of research into prompting techniques, including chain-of-thought (Wei et al. 2022), in-context learning (Dong et al. 2022), and retrieval augmented generation (Lewis et al. 2020), aimed at leveraging their generalization and reasoning capabilities for various downstream applications. In practice, providing suitable and descriptive (often long) prompts is essential for LLMs to generate useful responses for domain-specific tasks. However, longer prompts come at a significant inference cost for LLMs in terms of both time and finances (in the case of closed-source models). Therefore, striking a balance between the quality and the length of the prompt is a timely area of research in order to optimize inference performance vs. cost.
@@ 2 Related work
Prompt compression.
@@
Recent work on prompt compression aims to reduce the inference cost of LLM. A prominent direction in this area is model-agnostic prompt compression, which utilizes a pre-trained model to act as an information-based compressor. For instance, Kim et al. (2022) introduced a token pruning method in the forward pass of the LLM. In (Chen et al. 2023), a heuristic method was employed to recursively summarize the context. A limitation of these early works is that they require access to the pre-trained LLM as part of the compression method, which is not often feasible. Some of the more recent works developed solutions that do not require the LLM for compression. The most notable method in this direction is LLMLingua (Jiang et al. 2023b), which uses token-level perplexities to filter out semantically insignificant tokens that do not have high perplexity. Similar efforts have been performed by Selective Context (Li et al. 2023a), with decoder-only models (LLaMa (Touvron et al. 2023), GPT-2 (Radford et al. 2019)) for token-wise perplexity calculation and keep a pre-defined number of highest-perplexity tokens to achieve the required compression. LongLLMLingua (Jiang et al. 2023c) further developed this idea for question-aware context compression by adding document-level question relevance estimation prior to employing LLMLingua. In general, these methods suffer from the lack of capability to adapt to new domains.
@@
Another popular direction of prompt compression is trainable prompt compression methods, including soft prompt methods, sequence-to-sequence methods, and reinforcement learning methods, among others. Soft prompt methods directly pre-train or fine-tune a language model as the compressor (Wang et al. 2024), (Bulatov, Kuratov, and Burtsev 2022), (Chevalier et al. 2023), (Ge et al. 2024), and (Wang and Xiao 2024), which usually yield a high compression rate but have little interpretability or control over the compression rate. Sequence-to-sequence model to take the whole context as input and directly generate the compressed sentence (Xu, Shi, and Choi 2024). However, these methods have high latency due to the autoregressive nature of generation. Reinforcement learning for prompt compression, such as in (Laban et al. 2021), introduces a novel reward function to simplify complex text by jointly optimizing simplicity, fluency, salience, and guardrails. Another approach utilizes compression ratio as a reward function, conditioned on meeting a ROUGE metric threshold (Jung and Kim 2024), but may compromise on question-aware context compression tasks. Furthermore, reinforcement learning has been employed for token classification in efficient unsupervised sentence compression by fine-tuning pre-trained encoder (Ghalandari, Hokamp, and Ifrim 2022). Among more recent methods, Pan et al. (2024) trains a model that is designed to evaluate the information value of each specific lexical unit, which is then sorted by these values and pruned. However, the token-level compression results in a non-coherent sentence that hampers the semantics of the compressed prompt, resulting in a sub-optimal performance of the LLM. In this work, we focused on prompt compression by identifying and removing less relevant sentences from an input context.
Text embedding learning.
@@
The goal of text embedding learning is to generate text embeddings that capture the semantic meaning in a high-dimensional vector space, enabling various natural language processing tasks, such as text classification, clustering and similarity search. Earlier works in text embedding learning, such as GloVe (Pennington, Socher, and Manning 2014) and Word2Vec (Church 2017), focused on learning a word or token-level representation, while more recent works have explored sentence-level representation learning. For instance, Reimers and Gurevych (2019) proposed to fine-tune BERT-like architecture (Vaswani et al. 2017) for extraction of sentence embeddings that can be then utilized to measure text similarities. Later (Li et al. 2023b; Wang et al. 2022; Beltagy, Peters, and Cohan 2020) further improved the efficacy of the text embedding, especially for longer contexts. More recently, (BehnamGhader et al. 2024) utilized the vast knowledge of pre-trained LLMs to develop a strong sentence encoder. However, the sentence representation learned by such models is not context-aware, and no prior works focused on context-aware sentence encoding that is required for our prompt compression approach.
@@ 3.2 Dataset curation
To train the context-aware sentence encoder, we first create a dataset containing tuples of (long) contexts, questions, positive sentences, and negative sentences, called Context-aware Question-Relevance (CQR) Dataset. The ‘context’ is the text input that contains all the relevant information to answer the ‘question’, which is a meaningful query that asks some key information regarding the ‘context’. The ‘positives’ are defined as sentences within the ‘context’ that contain some, but not necessarily all, the relevant information to answer the question. Finally, the ‘negatives’ are ‘context’ sentences that contain no information relevant to answering the question. To generate a dataset of such tuples, we use a two-step approach where we first generate a set of questions, positives, and negatives, followed by a filtering step. We discuss these two steps below.
We start with the WikiText dataset (Merity et al. 2017) as the seed dataset for our synthetic data generation. This dataset consists of long Wikipedia pages, each containing factual information about historical/scientific concepts. First, we take a document from the original dataset which we consider as context $C$. This document consists of sentences $\{S_{i}\}_{i=1}^{K}$. Next, we sample ‘positive’ sentence $P$ from $S_{i}$. During the sampling procedure, we ensure that $P$ is a consistent English sentence that contains coherent information. We consider a sentence to be consistent and coherent if at least $\theta\%$ of its words are from the English dictionary and consist of ASCII-only characters. Then, we prompt a pre-trained upstream LLM $\psi$ to generate synthetic question-answer pairs $(Q_{j},A_{j})$ for this particular sentence $P$ while also considering its context $C$. This is performed using Prompt 1 below.
Prompt 1 (Question Prompt):
@@
Read the sentence in double brackets, namely, [[sentence]]
@@ 4.1 Datasets
We train our proposed context-aware sentence encoder on our own collected dataset (CQR dataset) described earlier. For evaluation, we follow the experimental setup of previous works (Jiang et al. 2023c) and report the performance of our proposed solution on LongBench (Bai et al. 2023) and ZeroSCROLLS (Shaham et al. 2023). LongBench is a dataset with different sub-tasks, including single-document QA, multi-document QA, and more. Similarly, ZeroSCROLLS is another dataset with different sub-tasks, including summarizing, QA, and few-shot learning.
@@
4.2 Implementation details
4.3 Evaluation protocols
We perform various studies on a number of downstream tasks based on specific commonly used evaluation criteria, which we present in the Appendix.
4.4 Results
Main results.
The main results of our proposed solution on the LongBench and ZeroSCROLLS datasets are presented in Table 2. Following prior works, we report the results on LongBench on the sub-tasks of a single document, multi-document, summarization, few-shot, synthetic, code, and the average over all the tasks. Following (Jiang et al. 2023c), the results are reported on 3,000 and 2,000 tokens, which is about 3$\times$ and 5$\times$ compression of the total token length. For ZeroSCROLLS, we present the performance with the same 3,000 and 2,000 tokens. As we observe in this table, our method outperforms the prior works by a considerable margin on LongBench with both 2,000 and 3,000 token constraints. More specifically, on 2,000 tokens, CPC outperforms others by 1.2 points on average and up to 8.3 points on individual tasks. On evaluation with 2,000 tokens, CPC shows a 1.5 points improvement on average. On ZeroSCROLLS, our proposed method also outperforms the previous state-of-the-art by 1.4 points on 3,000 tokens and 0.4 points on 2,000 tokens.
@@ 4.4 Results
Here, we discuss the latency of our proposed solution and compare it to prior works on prompt compression. We summarize the findings in Table 3. All the evaluations are conducted on the same A100 GPU used for training. The reported results are the average processing time of a sample of the LongBench. We report the results for our method in comparison to three prior SOTAs: LLMLingua (Jiang et al. 2023b), LLMLingua-2 (Pan et al. 2024), and LongLLMLingua (Jiang et al. 2023c). As we find from this study, our method is significantly faster than LLMLingua and LongLLMLingua, showing up to 27.5×27.5$\times$27.5 × and 10.93×10.93$\times$10.93 × average and median speedup. While LLMLingua-2 is slightly faster than our solution during inference, its performance is considerably worse than ours in all benchmarks and even worse than LongLLMLingua in most tasks. We also analyze the effect of the length of the original context on the execution latency of each of the above approaches. We depict these results in Figure 3. As we find from the Figure, the time complexity of our approach scales linearly with the growing length of the input context, while having relatively low latency compared to LongLLMLingua. On the other hand, LongLLMLingua has high latency for some of the samples with shorter context due to the dependence on the text structure, while our method is agnostic to the text structure.
@@
Next, we ablate the design choices of different components of the dataset collection procedure. We note that the dataset collection has several steps of verification and filtering: question and answer verification, similarity-based negative filtering, and KL-based negative filtering. In Table LABEL:tab:ablation_ds, we present an ablation study by removing each of these steps from our dataset collection pipeline. As we find from this table, removing question verification results in a 0.87 point drop in performance. This is caused by the fact that without question verification, the sentence and positive pairs are generated such that the question is directly fully answerable without the need to consider the context. Next, we find that removing the similarity-based negative filtering and KL-based filtering results in 1.04 and 1.67 points drop in performance, respectively, showing the importance of these steps in ensuring that negative does not contain any information regarding the question. the negatives to not contain any information regarding the question.
@@
We present a qualitative analysis of our prompt compression in Figure 4. Here, the first line shows a question, and the rest of the paragraph is the input context. The intensity of the colour indicates the relevance score of each sentence to the question predicted by our method. As we observe in this example, the sentences that are relevant to the question (‘What is the GhostVLAD approach?’) have higher scores than sentences that have less relevance.
@@ 5 Conclusion
In this work, we address the problem of prompt compression by developing a novel sentence-level context-aware compression method. To train our method, we first generated a new dataset of questions, positive, and negative pairs using an upstream LLM, followed by several steps of verification and filtering. We then use this dataset to contrastively train our novel context-aware sentence encoder. Detailed experiments demonstrate that our method is considerably faster than the state-of-the-art while showing considerable improvements on benchmark question-answering datasets. We believe our work will foster research in this promising direction of reducing the inference cost of LLMs.

### v3

@@ 1 Introduction
The advent of large language models (LLMs) has triggered a surge of research into prompting techniques, including chain-of-thought (Wei et al. 2022), in-context learning (Dong et al. 2022), and retrieval augmented generation (Lewis et al. 2020), aimed at leveraging their generalization and reasoning capabilities for various downstream applications. In practice, providing suitable and descriptive (often long) prompts is essential for LLMs to generate useful responses for domain-specific tasks. However, longer prompts come at a significant inference expense for LLMs in terms of both time and cost. Therefore, striking a balance between the quality and the length of the prompt is a timely area of research in order to optimize inference performance vs. cost.
@@ 2 Related work
In this section, we discuss the literature relevant to our work from two perspectives. First, we review previous works on prompt compression. Next, we provide an overview of the literature on sentence embedding learning, focusing on aspects relevant to our context-aware embedding module.
2.1 Prompt compression.
@@
Recent work on prompt compression aims to reduce the inference cost of LLMs. A prominent direction in this area is model-agnostic prompt compression, which utilizes a pre-trained model to act as an information-based compressor. For instance, Kim et al. (2022) introduced a token pruning method in the forward pass of the LLM. In Chen et al. (2023), a heuristic method was employed to recursively summarize the context. A limitation of these early works is that they require access to the pre-trained LLM as part of the compression method, which is not often feasible. Some of the more recent works developed solutions that do not require the LLM for compression. The most notable method in this direction is LLMLingua (Jiang et al. 2023b), which uses token-level perplexities to filter out semantically insignificant tokens that do not have high perplexity. Similar efforts have been performed by Selective-Context (Li et al. 2023a), with decoder-only models (LLaMa (Touvron et al. 2023), GPT-2 (Radford et al. 2019)) for token-wise perplexity calculation, which keeps a pre-defined number of highest-perplexity tokens to achieve the required compression. LongLLMLingua (Jiang et al. 2023c) further developed this idea for question-aware context compression by adding document-level question relevance estimation prior to employing LLMLingua. In general, these methods suffer from the lack of capability to adapt to new domains.
@@
Another common direction of prompt compression is trainable prompt compression methods, including soft prompt methods, sequence-to-sequence methods, and reinforcement learning methods, among others. Soft prompt methods directly pre-train or fine-tune a language model as the compressor (Wang et al. 2024; Bulatov, Kuratov, and Burtsev 2022; Chevalier et al. 2023; Ge et al. 2024; Wang and Xiao 2024), which usually yield a high compression rate but have little interpretability or control over the compression rate. Sequence-to-sequence models capture the entire context as input and directly generate the compressed sentence (Xu, Shi, and Choi 2024). However, these methods have high latency due to the autoregressive nature of generation. Reinforcement learning for prompt compression, such as in Laban et al. (2021), introduces a novel reward function to simplify complex text by jointly optimizing simplicity, fluency, salience, and guardrails. Another approach utilizes compression ratio as a reward function, conditioned on meeting a ROUGE metric threshold (Jung and Kim 2024), but may compromise on question-aware context compression tasks. Furthermore, reinforcement learning has been employed for token classification in efficient unsupervised sentence compression by fine-tuning pre-trained encoder (Ghalandari, Hokamp, and Ifrim 2022). Among more recent methods, Pan et al. (2024) trains a model that is designed to evaluate the information value of each specific lexical unit, which is then sorted by these values and pruned. However, the token-level compression results in a non-coherent sentence that compromises the semantics of the compressed prompt, resulting in a sub-optimal performance of the LLM. In this work, we focused on prompt compression by identifying and removing less relevant sentences from an input context.
2.2 Text embedding learning.
@@
The goal of text embedding learning is to generate text embeddings that capture the semantic meaning in a high-dimensional vector space, enabling various natural language processing tasks, such as text classification, clustering and similarity search. Earlier works in text embedding learning, such as GloVe (Pennington, Socher, and Manning 2014) and Word2Vec (Church 2017), focused on learning a word or token-level representation, while more recent works have explored sentence-level representation learning. For instance, Reimers and Gurevych (2019) proposed to fine-tune BERT-like architecture (Vaswani et al. 2017) for extraction of sentence embeddings that can be subsequently utilized to measure text similarities. Later Li et al. (2023b); Wang et al. (2022); Beltagy, Peters, and Cohan (2020) further improved the efficacy of the text embedding, especially for longer contexts. More recently, BehnamGhader et al. (2024) utilized the vast knowledge of pre-trained LLMs to develop a strong sentence encoder. However, the sentence representation learned by such models is not context-aware, and no prior works focused on context-aware sentence encoding that is required for our prompt compression approach.
@@ 3.2 Dataset curation
To train the context-aware sentence encoder, we first create a dataset containing tuples of (long) contexts, questions, positive sentences, and negative sentences, called Context-aware Question-Relevance (CQR) Dataset. The ‘context’ is the text input that contains all the relevant information to answer the ‘question’, which is a meaningful query that asks some key information regarding the ‘context’. The ‘positives’ are defined as sentences within the ‘context’ that contain some, but not necessarily all relevant information to answer the question. Finally, the ‘negatives’ are ‘context’ sentences that contain no information relevant to answering the question. To generate a dataset of such tuples, we use a two-step approach where we first generate a set of questions, positives, and negatives, followed by a filtering step. We discuss these two steps below.
We start with the WikiText dataset (Merity et al. 2017) as the seed dataset for our synthetic data generation. This dataset consists of long Wikipedia pages, each containing factual information about historical/scientific concepts. First, we take a document from the original dataset which we consider as context $C$. This document consists of sentences $\{S_{i}\}_{i=1}^{K}$. Next, we sample ‘positive’ sentence $P$ from $S_{i}$. During the sampling procedure, we ensure that $P$ is a consistent English sentence that contains coherent information. We consider a sentence to be consistent and coherent if at least $\theta\%$ of its words are from the English dictionary and consist of ASCII-only characters. Then, we prompt a pre-trained LLM $\psi$ to generate synthetic question-answer pairs $(Q_{j},A_{j})$ for this particular sentence $P$ while also considering its context $C$. This is performed using Prompt 1 below. Prompt 1 (Question Prompt):
@@
Read the sentence in double brackets, namely, [[sentence]].
@@ 4.1 Datasets
We train our proposed context-aware sentence encoder on our own collected dataset (CQR dataset) described earlier. For evaluation, we follow the experimental setup of previous works (Jiang et al. 2023c) and report the performance of our proposed solution on LongBench (Bai et al. 2023) and ZeroSCROLLS (Shaham et al. 2023). LongBench is a dataset with different sub-tasks, including single-document QA, multi-document QA, and more. Similarly, ZeroSCROLLS is another dataset with different sub-tasks, including summarization, QA, and few-shot learning. We also evaluate on domain-specific benchmarks such as medical paper summarization (PubMed (Cohan et al. 2018)), keyword extraction (Krapivin (Krapivin et al. 2009)), meeting transcript summarization (MeetingBank (Hu et al. 2023)) and TV show summarization (SummScreen (Chen et al. 2021)) datasets.
Method
Krapivin-2009
PubMed
MeetingBank
SummScreen
LLMLingua-2 36.7 26.24 16.73 21.57
LongLLMLingua 17.1 24.31 14.39 17.48
CPC 43.5 27.59 17.66 21.83
Table 3: Comparison to prior works on different domain-specific evaluation tasks.
Embed. Sin. Doc Mul. Doc Summ. FewShot Synth. Code AVG
LLMLingua-2 33.39 52.07 24.88 46.94 27.0 59.84 40.69
LongLLMLingua 26.70 39.25 22.66 58.45 13.75 27.73 31.42
CPC (Ours) 42.43 62.86 25.74 69.25 53.75 63.44 52.91
Table 4: Comparison to prior works with larger encoder.
4.2 Implementation details
4.3 Evaluation protocols
We perform various experiments on several downstream tasks based on commonly used protocols. We evaluate the effectiveness of our method on summarization problems by calculating the Rouge metric (Lin 2004) between the ground truth responses versus the model’s generated outputs given compressed prompts. For document QA tasks, we measure the F1 score between the response generated by the model and the ground truth response. For code completion tasks, we use the textual similarity metric based on the Levenshtein edit distance (Yujian and Bo 2007). For the different subsets of LongBench and ZeroSCROLLS, we follow the same evaluation protocol adopted by existing literature (Jiang et al. 2023c). For the keyword extraction, we evaluate the recall of the keywords that LLM extracts from a compressed context in relation to ground truth keywords.
@@ 4.3 Evaluation protocols
4.4 Results
Main results.
The main results of our proposed solution on the LongBench and ZeroSCROLLS datasets are presented in Table 2. Following prior works, we report the results on LongBench on the sub-tasks of a single document, multi-document, summarization, few-shot, synthetic, code, and the average over all the tasks. Following Jiang et al. (2023c), the results are reported on 3,000 and 2,000 tokens, which is about 3$\times$ and 5$\times$ compression of the total token length. For ZeroSCROLLS, we present the performance with the same 3,000 and 2,000 tokens. As we observe in this table, our method outperforms the prior works by a considerable margin on LongBench with both 2,000 and 3,000 token constraints. More specifically, on 2,000 tokens, CPC outperforms others by 1.2 points on average and up to 8.3 points on individual tasks. On evaluation with 2,000 tokens, CPC shows a 1.5 points improvement on average. On ZeroSCROLLS, our proposed method also outperforms the previous state-of-the-art by 1.4 points on 3,000 tokens and 0.4 points on 2,000 tokens.
@@ 4.4 Results
Domain generalization.
Since CPC effectively removes sentences that do not contain any useful information regarding the question or the prompt, it holds the potential for a wide range of tasks/domains. To this end, we investigate the performance of CPC on new domain-specific benchmarks described earlier in Section 4.1. The results of these experiments are presented in Table 3 with a 2,000-token constraint. As we observe from this table, CPC achieves significant improvements over prior methods across all domain-specific tasks. Specifically, CPC outperforms LLMLingua-2 by 6.8, 1.35, 0.93, and 0.26 points on Krapivin-2009, PubMed, MeetingBank, and SummScreen, respectively.
Larger backbone.
To evaluate the scalability of our method with a larger model, we compare the performance of CPC against prior works with a larger LLM, GPT-4o, as the backbone. The results of this study are presented in Table 4. As we find from this table, CPC consistently outperforms existing methods across all subtasks of the LongBench. On average, CPC outperforms LLMLingua-2 and LongLLMLingua by 12.22 and 21.49 points, respectively.
@@
Here, we discuss the latency of our proposed solution and compare it to prior works on prompt compression. We summarize the findings in Table 5. All the evaluations are conducted on the same A100 GPU used for training. The reported results are the average processing time of a sample of the LongBench. We report the results for our method in comparison to three prior SOTAs: LLMLingua (Jiang et al. 2023b), LLMLingua-2 (Pan et al. 2024), and LongLLMLingua (Jiang et al. 2023c). As we find from this study, our method is significantly faster than LLMLingua and LongLLMLingua, showing up to 27.5×27.5$\times$27.5 × and 10.93×10.93$\times$10.93 × average and median speedup. While LLMLingua-2 is slightly faster than our solution during inference, its performance is considerably worse than ours in all benchmarks and even worse than LongLLMLingua in most tasks. We also analyze the effect of the length of the original context on the latency of each of the above approaches. We depict these results in Figure 3. As we find from the Figure, the time complexity of our approach scales linearly with the growing length of the input context, while having relatively low latency compared to LongLLMLingua. On the other hand, LongLLMLingua has high latency for some of the samples with shorter context due to the dependence on the text structure, while our method is agnostic to the text structure.
@@
Next, we ablate the design choices of different components of the dataset collection procedure. We note that the dataset collection has several steps of verification and filtering: question and answer verification, similarity-based negative filtering, and KL-based negative filtering. In Table LABEL:tab:ablation_ds, we present an ablation study by removing each of these steps from our dataset collection pipeline. As we find from this table, removing question verification results in a 0.87 point drop in performance. This is caused by the fact that without question verification, the sentence and positive pairs are generated such that the question is directly fully answerable without the need to consider the context. Next, we find that removing the similarity-based negative filtering and KL-based filtering results in a 1.04 and 1.67 points drop in performance, respectively, showing the importance of these steps in ensuring that the negative does not contain any information regarding the question.
@@
We present a qualitative analysis of our prompt compression method in Figure 4. Here, the first line shows a question, and the rest of the paragraph is the input context. The intensity of the colour indicates the relevance score of each sentence to the question predicted by our method. As we observe in this example, the sentences that are relevant to the question (‘What is the GhostVLAD approach?’) have higher scores than sentences that have less relevance.
@@ 5 Conclusion
In this work, we address the problem of prompt compression by developing a novel sentence-level context-aware compression method. To train our method, we first generated a new dataset of questions, positive, and negative pairs using a pre-trained LLM, followed by several steps of verification and filtering. We then use this dataset to contrastively train our novel context-aware sentence encoder. Detailed experiments demonstrate that our method is considerably faster than the state-of-the-art while showing considerable improvements on benchmark question-answering datasets. Additionally, our method shows strong domain generalization and robust performance with larger encoders. We believe our work will foster research in this promising direction of reducing the inference cost of LLMs.